009 10.0 points A racing car has a mass of 1810 kg. The acceleration of gravity is 9.8 m/ What is its kinetic energy if it has a speed of 131 km/h? Answer in units of J. 010 10.0 points Calculate the speed of an 7.4 x 101kg airliner with a kinetic energy of 1.0 x 10J. Answer in units of m/s.
Answer #1

1) Mass = 1810 kg

Speed = 131 km/h = 131 x (5/18) = 36.4 m/s

Then Kinetic Energy = \frac{1}{2}mv^2 = \frac{1}{2}(1810)(36.4)^2 = 1199088.8J \simeq 1200kJ

2) Given mass = 7.4 x 104 kg

Kinetic Energy = 1 x 109 J

Then K.E = \frac{1}{2}mv^2

V) 7.4×101 164.4m/s